OpenStudy (anonymous):

-3x-y=8 7x+y=-12 linear system using elimination .

OpenStudy (anonymous):

-3x-y=8 7x+y=-12 --------- So you have to solve this by elimination. Can you sum them?

OpenStudy (anonymous):

4x =-4 which becomes -1x then -3(-1)-y=8 Which is ? 5?

OpenStudy (anonymous):

So you found x = -1 and did -3(-1)-y=8, which is the same as 3-y=8 => -y = 8 - 3 => -y=5. It gives us y=-5.

OpenStudy (anonymous):

Thanks . So when x is negative , the y would become negative also ?

OpenStudy (anonymous):

It depends on the equation. Sometimes yes, sometimes nope.

OpenStudy (anonymous):

Oh okay.